Colton Enterprises experienced the following events for Year 1, the first year of operation:

  1. Acquired $50,000 cash from the issue of common stock.
  2. Paid $13,500 cash in advance for rent. The payment was for the period April 1, Year 1, to March 31, Year 2.
  3. Performed services for customers on account for $102,000.
  4. Incurred operating expenses on account of $42,500.
  5. Collected $78,000 cash from accounts receivable.
  6. Paid $36,000 cash for salary expense.
  7. Paid $34,000 cash as a partial payment on accounts payable.

Adjusting Entries

  1. Made the adjusting entry for the expired rent. (See Event 2.)
  2. Recorded $5,400 of accrued salaries at the end of Year 1.

Events for Year 2

  1. Paid $5,400 cash for the salaries accrued at the end of the prior accounting period.
  2. Performed services for cash of $51,000.
  3. Purchased $4,300 of supplies on account.
  4. Paid $15,000 cash in advance for rent. The payment was for one year beginning April 1, Year 2.
  5. Performed services for customers on account for $118,000.
  6. Incurred operating expenses on account of $56,500.
  7. Collected $104,000 cash from accounts receivable.
  8. Paid $54,000 cash as a partial payment on accounts payable.
  9. Paid $33,000 cash for salary expense.
  10. Paid a $12,000 cash dividend to stockholders.

Adjusting Entries

  1. Made the adjusting entry for the expired rent. (Hint: Part of the rent was paid in Year 1.)
  2. Recorded supplies expense. A physical count showed that $650 of supplies were still on hand.

c. Prepare a trial balance for Year 1.
Answer #1
Colton Enterprises
Trial Balance
Year 1
Debit Credit
Cash $44500
Accounts receivable 24000
Prepaid rent 3375
Salaries payable 5400
Accounts payable 8500
Service revenue 102000
Common stock 50000
Operating expenses 42500
Salaries expense 41400
Rent expense 10125
Total $165900 $165900

Cash= $50000-13500+78000-36000-34000= $44500

Accounts receivable= $102000-78000= $24000

Prepaid rent= $13500-(13500*9/12)= $3375

Accounts payable= $42500-34000= $8500

Salaries expense= $36000+5400= $41400

Rent expense= $13500*9/12= $10125